OpenStudy (anonymous):

What is the simplified form of 2 over x squared plus x minus 1 over x

OpenStudy (anonymous):



OpenStudy (anonymous):

create a common denominator... can you factor the left denominator

OpenStudy (anonymous):

This is new material, so I don't really understand

OpenStudy (anonymous):

alright so when you add fractions you need the denominators to be the same correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{1}{4}\] cannot be added unless the first part has 4 in the denominator. So you need to look and see that 1/2 is missing a multiple of 2 in the denominator (2*2=4). If you multiply by one in the form (2/2), it will not change the fraction but it will change the denominator. \[\frac{2}{2}\frac{1}{2}+\frac{1}{4}\] \[\frac{2}{4}+\frac{1}{4}=\frac{3}{4}\]

OpenStudy (anonymous):

this is essentially the same with variables or functions also so you must find what the second part is missing in order to subtract them

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

Thank you nevermind I got the answerr

OpenStudy (anonymous):

no problem